OpenStudy (anonymous):

How I can find this value 0.45=1-exp((-a/2))

OpenStudy (ash2326):

We have \[0.45= 1-e^{\frac{-a}{2}}\] could you simplify it a bit?

OpenStudy (anonymous):

\[\frac{ 9 }{ 20 }=1-\frac{ 1 }{ \sqrt{a} }\] like this ?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{e^a} }\]

OpenStudy (ash2326):

I think we have like this \[e^{\frac{-a}{2}}=0.55\] Do you follow?

OpenStudy (anonymous):

yes , i guess

OpenStudy (anonymous):

how did you 0,55 ?

OpenStudy (ash2326):

we had \[0.45=1-e^{\frac{-a}{2}}\] subtract both sides by 1 \[0.45-1=1-e^{\frac{-a}{2}}-1\] now we get \[-0.55=\cancel {1}-e^{\frac{-a}{2}}-\cancel {1}\] \[-0.55=-e^{\frac{-a}{2}}\] Multiply both sides by -1 and swap \[e^{\frac{-a}{2}}=0.55\] Do you follow?

OpenStudy (anonymous):

i see

OpenStudy (ash2326):

do you know logarithm?

OpenStudy (anonymous):

used to know it but i dont recall it

OpenStudy (ash2326):

okay natural log or log to the base e is defined as \[\log_{e} x=y\] \[\text{this implies}\\ x= e^{y}\] do you follow?

OpenStudy (anonymous):

yes i follow

OpenStudy (ash2326):

okay let's apply log to the both sides in our expression we have \[e^\frac{-a}{2}=0.55\] applying log to the base e \[\log_e e^\frac{-a}{2}=\log_e 0.55\] now per logarithmic property \[\log_a a^b=b\] so we get \[\frac{-a}{2}=\log_e 0.55\] Do you follow these steps?

OpenStudy (anonymous):

k

OpenStudy (ash2326):

now to find value of \[\log_e 0.55\] we need to refer log tables, do yo have one? or is this value given in question..

OpenStudy (anonymous):

i don't a table but log 0.55 is like a negative numbers

OpenStudy (anonymous):

log (0.55) =-.519274621

OpenStudy (ash2326):

yes, it will be negative as 0.55 is less than e We can use a calculator but it that allowed in your homework?

OpenStudy (anonymous):

im allow to use a program call maple

OpenStudy (anonymous):

but i still need to find the value of a

OpenStudy (ash2326):

check it if there is log function there..

OpenStudy (anonymous):

what i need to check?

OpenStudy (ash2326):

please check this link.. http://www.maplesoft.com/support/help/maple/view.aspx?path=ln

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-.5978370008/ln(e)

OpenStudy (ash2326):

okay so we have \[-a/2=-0.5978\] I think you can find a now...

OpenStudy (anonymous):

is 1.195674002

OpenStudy (ash2326):

yes, just simple math

OpenStudy (anonymous):

k thx so you much for the help

OpenStudy (ash2326):

No problem.